OpenStudy (anonymous):

express as a single logarithm: 3logb(3(sqrt(x))-2logb(x)

OpenStudy (anonymous):

\[3\log(3\sqrt3)-2\log(x)\] like that ?

OpenStudy (anonymous):

no that is not right, how about \[3\log(3\sqrt{x})-2\log(x)\]

OpenStudy (anonymous):

start by turning the coefficients in to exponents and write \[\log((3\sqrt{x})^3)-\log(x^2)\]

OpenStudy (anonymous):

then the subtraction becomes a division, and you get \[\log\left(\frac{27(\sqrt{x})^3}{x^2}\right)\]

OpenStudy (anonymous):

this part \(\frac{27(\sqrt{x})^3}{x^2}\) needs some cleaning up, but that is the idea